OpenStudy (anonymous):

Determine the value of A in the rational equation given. (A/(x-4)) + (1/(x-5)) = (-x+6)/(x^2 = 9x + 20)

OpenStudy (anonymous):

@geerky42 Mind helping me?

myininaya (myininaya):

Too many equal signs.

geerky42 (geerky42):

\[\dfrac{A}{x-4} + \dfrac{1}{x-5} = \dfrac{-x+6}{x^2 - 9x + 20}\]Right?

myininaya (myininaya):

I bet your right @geerky42 since (x-4)(x-5) is x^2-9x+20

OpenStudy (anonymous):

@geerky42 sorry I was brb

OpenStudy (anonymous):

yes

geerky42 (geerky42):

Can you combine fractions on left side?

geerky42 (geerky42):

\[\dfrac{A}{x-4} + \dfrac{1}{x-5} = ~?\]

OpenStudy (anonymous):

I simplified and got xA-5A+2/(x-4)(x-5) but i'm sure you aren't suppose to simplify